The same problem under FIFO method

Muskoge Company uses a process-costing system. The
Answer #1

6.37: Molding department:

Molding Department
Units, BWIP 10000
Units Started at February 25000
Total units to account for 35000
Units, BWIP 10000
Units Started and completed 20000
Units, EWIP 5000
Total units accounted for 35000
Direct Materials Conversion Costs
United started and completed 20000 20000
Units, BWIP 0 6000
Units, EWIP 5000 4000
Total Equivalent units 25000 30000

Calculations and explanations: BWIP units = BWIP units*(1-beginning inventory % completion) = 10,000*(1-40%)

Cost computation:

Direct Material Conversion costs
Cost in BWIP 22,000 13,800
Cost added 56,250 103,500
Total cost 78,250 117,300
Total Units
Unit Costs 2.25 3.45 5.7

Unit costs = costs added/total equivalent units

Cost, BWIP 35,800
Cost Added 159,750
Total Cost to account for 195,550

Calculations: costs added = direct materials+conversion costs = 56250+103500 = 159750

Cost of Goods Started and transferred out 114000
Incurred cost, BWIP 20700
Cost, BWIP 35,800
Total cost of goods transferred out     170,500
Cost, EWIP
Direct Materials 11250
Conversion Costs 13800 25050
Total cost accounted for     195,550

Calculations: costs of goods started and transferred out = total units costs*units started and completed

Incurred cost, BWIP = conversion unit costs*units BWIP

Cost, BWIP = direct materials+conversion costs

Direct materials = Total unit costs of direct materials*Units EWIP (Direct materials)

Conversion costs = Total unit costs of conversion*Units EWIP (conversion)

Journal entries:

Dr Cr
WIP - molding 56,250
Materials Inventory 56,250
WIP-Molding 103,500
Conversion Costs 103,500
WIP-Assembly 170,500
WIP-Molding 170,500

The same formulas and method has been used for Assembly department.

Assembly
Units, BWIP 8,000
Units Started at February 30,000
Total units to account for 38,000
Units, BWIP 8,000
Units Started and completed 27,000
Units, EWIP 3,000
Total units accounted for 38,000
Direct Materials Conversion Costs Transferred in
United started and completed 27,000 27,000 27,000
Units, BWIP 8,000 4,000 0
Units, EWIP 0 1500 3,000
Total Equivalent units 35,000 32,500 30,000
Direct Material Conversion costs Transferred in
Cost in BWIP 0 16,800 45,200
Cost added 39,550 136,500 170500
Total cost 39,550 153,300 215,700 total cost
Unit Costs 1.13 4.2 5.683333 11.01333
Cost, BWIP 62,000
Cost Added 346,550
Total Cost to account for 408,550
Cost of Goods Started and transferred out 297360
Incurred cost, BWIP 25840
Cost, BWIP 62,000
Total cost of goods transferred out 385,200
Cost, EWIP
Direct Materials 6300
Conversion Costs 17050 23350
Total cost accounted for 408,550

Journal entries:

Dr Cr
WIP - Assembly 39,550
Materials Inventory 39,550
WIP - Assembly 136,500
Conversion Costs 136,500
Finished Goods 385,200
WIP - Assembly 385,200